# Video Is Worth a Thousand Images: Exploring the Latest Trends in Long Video Generation

### v1

@@ 7. Conclusion & Future Trends
One of the significant open research areas in long video generation is the generation of longer videos that accurately reflect spatial, temporal, and physical dynamics. A key challenge is the need for large-scale video datasets with comprehensive spatial, temporal, and physical context (e.g., trajectories, shadows, interactions). Existing large-scale datasets like HD-VG-130M  [127] offer scale but have limitations in terms of caption quality (e.g., captions are restricted to 15-20 words and lack rich spatial and temporal contextual information). On the other hand, datasets such as VideoInstruct-100K  [64] provide rich spatial and temporal context but fall short in scale. Developing datasets that balance both scale and rich context is critical for advancing long video generation research. In addition to datasets, measuring generated video quality presents another challenge. Current state-of-the-art metrics, such as FETV  [59], MiraBench  [42], and VBench [39], rely on manual human feedback to assess video quality, which is time-consuming, subjective, and challenging to scale. Future research should focus on developing fully automated metrics that can evaluate the quality of generated videos in a more scalable and objective manner.
Another open area of research in long video generation is the integration of audio. Currently, most commercial video generation models, such as SORA and Stability AI, do not produce accompanying audio. Developing methods to generate audio that aligns seamlessly with visual content is crucial for creating immersive and comprehensive videos, making this an important focus in the field of long video generation.
Long video generation promises to revolutionize multiple fields, including entertainment, education, virtual reality, and game development. However, it also introduces significant challenges, such as the potential for fake video creation, bias, violence, and moral concerns. Additionally, issues like hallucinations can limit the applicability of generative videos, particularly in domains like education and science. In conclusion, this survey offers readers an in-depth overview of the state of the art in long video generation, highlighting key research areas and opportunities for future exploration.

### v2

@@ 8. Conclusion & Future Trends
One of the significant open research areas in long video generation is the generation of longer videos that accurately reflect spatial, temporal, and physical dynamics. A key challenge is the need for large-scale video datasets with comprehensive spatial, temporal, and physical context (e.g., trajectories, shadows, interactions). Existing large-scale datasets, such as HD-VG-130M  [142], offer scale but have limitations in terms of caption quality (e.g., captions are restricted to 15-20 words and lack rich spatial and temporal contextual information). On the other hand, datasets such as VideoInstruct-100K  [137] provide rich spatial and temporal context but fall short in scale. Developing datasets that balance both scale and rich context is critical for advancing long video generation research. In addition to datasets, measuring the quality of generated videos presents another challenge. Current state-of-the-art metrics, such as FETV  [156], MiraBench  [139], and VBench [157], rely on manual human feedback to assess video quality, which is time-consuming, subjective, and challenging to scale. Future research should focus on developing fully automated metrics that can objectively evaluate the quality of generated videos in a more scalable manner.
Another open area of research in long video generation is the integration of audio. Currently, most commercial video generation models, such as SORA and Stability AI, do not produce accompanying audio. Developing methods to generate audio that aligns seamlessly with visual content is crucial for creating immersive and comprehensive videos, making this a key focus in the field of long-form video generation.
Long video generation promises to revolutionize multiple fields, including entertainment, education, virtual reality, and game development. However, it also introduces significant challenges, such as the potential for fake video creation, bias, violence, and moral concerns. Additionally, issues like hallucinations can limit the applicability of generative videos, particularly in domains like education and science. In conclusion, this survey provides readers with an in-depth overview of the current state of the art in long video generation, highlighting key research areas and opportunities for future exploration.